# Guaranteed confidence-band enclosures for PDE surrogates

### v1

@@ 3.1 Fitting zonotopic confidence sets
Figure 4: Illustration of fitting a zonotope using the principal components.
A singular value decomposition (SVD) of the data $X=U\Sigma V^{\top}$ yields a diagonal matrix $\Sigma$ of (decreasing) singular values and $V^{\top}$ matrix of singular (or eigen-) vectors. The singular values provide a natural ranking of the contribution of each eigenvector on the variance of $X$. An enclosing hyperrectangle can be easily found by computing the data $\min$ and $\max$ over each dimension of $U$. Upon converting this hyperrectangle (more details appendix B.1) to a zonotope representation, the zonotope $\mathcal{Z}_{U}=\langle c_{U},G_{U}\rangle$ can be projected back to the original space by $\mathcal{Z}_{X}=\langle V\Sigma c_{U},V\Sigma G_{U}\rangle$. This yields a zonotope whose generators are aligned with the eigenvectors of the dataset. Figure 4 illustrates this on a half-moon dataset. Although simple, quick to compute, and scalable, an obvious downside to this method is that it only yields zonotopes with generators as the same number of dimensions of $X$.
@@
A fast and simple, but potentially inaccurate, method is to take $p_{Z}$ to be the data point $X_{i}$ with the greatest Euclidean depth w.r.t the sample mean $\mu$
@@
We find this method also fast, scalable, and more accurate than the Euclidean depth when the data covariance can be inverted. (17) use a metric similar to $d_{\text{M}}$ as a non-conformity score, the level-sets of which are elliptical.
@@
3.2 Calibrating zonotopic confidence sets
@@
For mostly illustrative purposes, we begin by showing how $Z^{\alpha}_{p_{Z}}$ can be calibrated from a known multivariate distribution, whose cdf $F_{X}$ or density $f_{X}$ are available. This could perhaps be useful for some applications, but the next method (calibrating from sample data) is likely to find wider use. Given a zonotope $\mathcal{Z}$, ideally containing the range of $f_{X}$ or otherwise some large probability region ($\mathbb{P}_{X}(\mathcal{Z})\approx 1$) if $X$ is unbounded, and a $p_{Z}$ ideally near the mode, the structure can be calibrated by integrating the density $f_{X}$ in the sets $\mathcal{Z}^{\alpha}_{p_{Z}}$:
@@
We note that computing the exact membership score $\alpha_{i}=\sup\{\alpha\in[0,1]\;\mid\;X_{i}\in\mathcal{Z}^{\alpha}_{p_{Z}}\}$ may be challenging, as the set must be varied with $\alpha$ until $X_{i}\notin\mathcal{Z}^{\alpha}_{p_{Z}}$, which could be relatively well performed using mathematical optimisation for certain set representations. We however suggest a simple method, where $\alpha$ uniformly discretised in a grid in $A=\{0,0.111,\dots,0.999,1\}$, and computing
@@
Although the exact supremum isn’t found, this still gives conservative results, as two collection of scores $\alpha_{i}\leq\alpha_{j}$ will yield a lower bound on probabilistic bound on 3. I.e. larger confidence set $\mathcal{Z}_{p_{Z}}^{\alpha_{i}}\supseteq\mathcal{Z}_{p_{Z}}^{\alpha_{j}}$. We suggested that the discretisation of the $\alpha$ values is as least a large as the data set, to avoid multiple repetitions of the same $\alpha_{i}$ values as much as possible. Also note that like in conformal prediction, the maximum number of unique prediction sets that can be obtained is the number of calibration data points provided. Figure 6 gives three examples of calibrating $\mathcal{Z}^{\alpha}_{p_{Z}}$ using this method with different data lengths.
@@
Some potential alternative frameworks exist for reliably calibrating $\mathcal{Z}^{\alpha}_{p_{Z}}$, including inferential models (15), confidence-structures (3), frequency-calibrated belief functions (8), and scenario theory (7). There are likely strong links between these methods and what is proposed here.
4 Functional confidence-bands
Given a pre-trained model $\hat{f}:X\mapsto F$, which maps from Euclidean space $X\in\mathbb{R}^{m}$ (space of PDE inputs) to a space of functions $F\in\mathcal{F}$ (PDE solutions), discretised on a fine grid $F_{1}=[F_{1}(y_{1}),\dots,F_{1}(y_{l})]\in\mathbb{R}^{l}$, we can compute the error of $\hat{f}$’s error on some unseen data $\{(X_{1},F_{1}),\dots,(X_{n},F_{n})\}$ with $e_{i}=F_{i}-\hat{f}(X_{i})$. We may then reduce the dimension of $e_{i}$ using an SVD: $e=U\Sigma V^{\top}$, where the singular values indicate the variance contribution of each eigenvector to the variance of $e_{i}$. We may truncate the dimensions of the data, capturing some high (often $99\%$) of the overall variance, and projecting the data $e_{i}$ onto these remaining dimensions, a low dimension representation of $\hat{f}$’s error: $U_{i}$. Depending on the size of these dimensions, one of the above fitting methods may be applied to find $\mathcal{Z}^{\alpha}_{p_{Z}}$. We find the ‘rotated hyperrectangle’ method and the Mahalanobis depth scale well to high dimensions, with the convex hull method being tighter for smaller dimensions. $\mathcal{Z}^{\alpha}_{p_{Z}}$ may then be calibrated, giving $\mathbb{P}(U_{n+1}\in\mathcal{Z}^{\alpha})\geq 1-\alpha$. Mapping $\mathcal{Z}^{\alpha}_{p_{Z}}$ through the linear transformation back to $\mathbb{R}^{l}$ is straightforward for the underlying zonotope, and for the probabilistic part $\alpha$, propagate each zonotope level-set independently $\mathcal{R}^{\alpha}=f(\mathcal{Z}^{\alpha}_{p_{Z}})$, i.e. perform one transformation for each confidence-level, with the $\alpha$-guaranteed being retained by the set. Note we may perform this since we compute a multivariate confidence set, this would not be the case had each dimension been fit independently.
These nested sets can be added to the prediction of $\hat{f}$, yielding the desired functional prediction set
$\mathbb{C}^{\alpha}_{F}=\hat{f}(X_{n+1})+\mathcal{R}^{\alpha}.$
@@ 4.1 Bounding truncation error
By taking a bounding box of $F_{i}$’s projection onto the truncated eigenvectors, one can obtain an estimate of the range of uncertainty in these dimensions, that is, the range of $\mathcal{Z}$ had we included these dimensions. We capture this error with a simple bounding hyperrectangle $E$ found using the data minimum and maximum in each dimension. Upon taking the Cartesian product $\mathcal{R}^{\alpha}=\mathcal{Z}^{\alpha}_{p_{Z}}\times E$, one obtains a zonotope in high dimensions which contracts in the important directions as $\alpha$ varies, but remains constant in the truncated dimensions. The $\alpha$-guarantee remains unchanged due to this operation because: 1) the number of data points remains unchanged (only their dimension), and 2) the data is totally bounded (and remains so as $\alpha$ changes) in the extra dimensions. Indeed, there is a slight performance-loss when performing this (our probabilistic bound becomes looser), however this is quite minor as these extra dimensions don’t affect $F_{i}$ dramatically. The Cartesian product between two zonotopes $\mathcal{Z}_{X}\subset\mathbb{R}^{n}$ and $\mathcal{Z}_{Y}\subset\mathbb{R}^{m}$ produces an additional zonotope $\mathcal{Z}_{X\times Y}\subset\mathbb{R}^{n\times m}$, whose centre and generator are a simple concatenation of the centres and generators of $\mathcal{Z}_{X}$ and $\mathcal{Z}_{Y}$: $c_{X\times Y}=(c_{X}||c_{Y})$ and $G_{X\times Y}=(G_{X}||G_{Y})$.
4.2 PDE surrogate example
We use LazySets.jl from (10) for the set construction, and NeuralOperators.jl for the base Sci-ML model. Otherwise, a supplementary code repository will be released for those wishing to reproduce this papers results.
Figure 7: Examples of constructed confidence-bands of the FNO for Burger’s equation, predicted on test data set. The ground truth is shown in yellow.
@@ 4.2 PDE surrogate example
To demonstrate the methodology, we construct functional confidence bands on a Fourier neural operator (FNO) for the Burger’s equation, trained on data supplementary data provided from (14), which was built by sampling a numerical PDE solver. The FNO model was trained on an A100 GPU using NeuralOperators.jl, using $2048$ points to train the underlying model, $1000$ point for calibrating the confidence bands, and $1048$. Additional information about the training setup, and about the underlying PDEs, can be found in appendix C. The predictions from the FNO are discretise onto a grid of 1024 points. A $32$-dimensional rotated hyperrectangle was fitted on the important directions the error’s SVD, with $p_{Z}$ found using the Mahalanobis depth. The entire calibration process took $75$s on a modern laptop, with most of the computation spent on computing the membership 4 of the calibration data. Once calibrated, computing a prediction is timed at $0.02$s. Four realisations of the training dataset are shown in Figure 7, with the ground truth shown in yellow.
@@
We also note that a 1024 nested zonotopic set is predicted for the output of the FNO, what’s shown in Figure 7 is the axis projection of this set. If we inspect the individual axis, the dependence of the field can also be seen captured.
5 Relationships to other ideas
5.1 Non-convex confidence sets
It is fairly straightforward to extend the above to other set-representations, for example Taylor-models or polynomial zonotopes, with examples shown in Figure 8. However, a compact representation, like our proposition 1 is challenging for more complex set representations. The below sets were constructed by first constructing a $\mathcal{Z}^{\alpha}_{p_{Z}}\subseteq[-1,1]^{p}$, and mapping this through the corresponding Taylor polynomial. Although this allows us to compute level-sets (and therefore memberships 4), an efficient method to fit these sets to data would need to be proposed.
Figure 8: Example nested polynomial zonotopes.
This more sophisticated set-representation could allow us to perform a more aggressive non-linear dimension reduction. A potentially interesting application of this work would be to probabilistically fit from data the set-representations used in dynamical system verification.
5.2 Relationship to conformal prediction
@@
The proposed method is similar to conformal prediction, but where we directly solve for a valid prediction set, instead of taking a level-set of a non-conformity scoring function. Indeed, the membership function of $\mathcal{Z}^{\alpha}_{p_{Z}}$ is the non-conformity score in conformal prediction. If one could find an efficient (perhaps analytical) expression for this membership function, in terms of $\alpha$ and the underlying set-representation, then the calibration procedure would be made much more efficient (corresponding to a simple function evaluation), without the need to check the set-membership of $\mathcal{Z}^{\alpha}_{p_{Z}}$ in 4, which is a costly part of the proposition. This would also additionally greatly simplify the implementation.
@@
We therefore share many of the advantages and dis-advantages of conformal prediction, including finite-sample guarantees and being model agnostic. Although we show example of regression problems, we believe our method could apply to multivariate classification problems.
6 Limitations
@@
Our method requires the same assumption as conformal prediction, namely exchangeability, we therefore suffer similar limitations. The guarantee is lost if the dataset is not exchangeable (for example if it changes over time). We also give only marginal coverage (averaged for all samples of $X_{i}$), rather than the stronger conditional coverage (conditioned for particular $x$), which is more desirable for surrogate modelling. Finally, we mention that our method would greatly benefit for refined methods for fitting set-representations to data. Although the routines we proposed are simple, we do not doubt that more sophisticated methods can be found, to improve the tightness of the fitted set.
Acknowledgements

### v2

@@ 3.1 Fitting zonotope prediction sets
Figure 4: Illustration of fitting a zonotope using the principal component analysis. The data is rotated into the PCA basis, a bounding hyperrectangle is fit in this space, then mapped back to the original coordinates. The resulting zonotope has generators aligned with the principal components, capturing the dominant directions of variation.
A singular value decomposition (SVD) of the data $X=U\Sigma V^{\top}$ yields a diagonal matrix $\Sigma$ of (decreasing) singular values and $V^{\top}$, the matrix of right singular vectors. The singular values provide a natural ranking of each eigenvector’s contribution to the variance of $X$. An enclosing hyperrectangle can be easily found by computing the data $\min$ and $\max$ over each dimension of $U$. Upon converting this hyperrectangle (more details appendix B.1) to a zonotope representation, the zonotope $\mathcal{Z}_{U}=\langle c_{U},G_{U}\rangle$ can be projected back to the original space by $\mathcal{Z}_{X}=\langle V\Sigma c_{U},V\Sigma G_{U}\rangle$. This yields a zonotope whose generators are aligned with the eigenvectors of the dataset. Figure 4 illustrates this on a half-moon dataset. Although simple, quick to compute, and scalable, an obvious downside to this method is that it only yields zonotopes with generators as the same number of dimensions of $X$.
@@
A fast, but potentially inaccurate, method is to take $p_{Z}$ to be the data point $X_{i}$ with the greatest Euclidean depth w.r.t the sample mean $\mu$
@@
We find this method also fast, scalable, and more accurate than the Euclidean depth when the data covariance can be inverted. Messoudi et al. [2022] use a metric similar to $d_{\text{M}}$ as a non-conformity score, the level-sets of which are elliptical.
@@
3.2 Calibrating zonotope prediction sets
@@
For mostly illustrative purposes, we begin by showing how $Z^{\alpha}_{p_{Z}}$ can be calibrated from a known multivariate distribution, whose cdf $F_{X}$ or density $f_{X}$ are available. This could perhaps be useful for some applications, but the next method (calibrating from sample data) is likely to find wider use. Given a zonotope $\mathcal{Z}$ (ideally containing the range of $f_{X}$ or otherwise some large probability region ($\mathbb{P}_{X}(\mathcal{Z})\approx 1$ if $X$ is unbounded) and a $p_{Z}$ ideally near the mode, the structure can be calibrated by integrating the density $f_{X}$ in the sets $\mathcal{Z}^{\alpha}_{p_{Z}}$:
@@
Figure 6: Effect of grid size in membership computation (equation 4). The example computes the prediction sets of a 2-dimensional standard normal gaussian. A large 2M calibration points and 200M points for empirical coverage $\mathbb{P}(\mathcal{Z}^{\alpha})$ are used to isolate the effect of the discretisation.
We note that computing the exact membership score $\alpha_{i}=\sup\{\alpha\in[0,1]\;\mid\;X_{i}\in\mathcal{Z}^{\alpha}_{p_{Z}}\}$ may be challenging, as the set must be varied with $\alpha$ until $X_{i}\notin\mathcal{Z}^{\alpha}_{p_{Z}}$, which could be relatively well performed using mathematical optimisation for certain set representations. We however suggest a simple method, where $\alpha$ is uniformly discretised in a grid in $A=\{0,0.111,\dots,0.999,1\}$, and computing
@@
Although the exact supremum isn’t found, this still gives conservative results, as two collection of scores $\alpha_{i}\leq\alpha_{j}$ will yield a lower bound on probabilistic bound on 3. I.e. larger prediction set $\mathcal{Z}_{p_{Z}}^{\alpha_{i}}\supseteq\mathcal{Z}_{p_{Z}}^{\alpha_{j}}$. This claim is evidenced with a numerical experiment in Figure 6, where one can observe that irrespective of the discretisation a guarantee can be obtained, however the “resolution” of the sets is effected: the more grid points used, the tighter the bounds are. We suggested that the discretisation of the $\alpha$ values is as least a large as the data set, to avoid multiple repetitions of the same $\alpha_{i}$ values as much as possible. Also note that like in conformal prediction, the maximum number of unique prediction sets that can be obtained is the number of calibration data points provided. Figure 7 gives three examples of calibrating $\mathcal{Z}^{\alpha}_{p_{Z}}$ using this method with different data lengths.
@@
Some potential alternative frameworks exist for reliably calibrating $\mathcal{Z}^{\alpha}_{p_{Z}}$, including inferential models [Martin and Liu, 2015], confidence-structures [Balch, 2012], frequency-calibrated belief functions [Denœux and Li, 2018], and scenario theory [De Angelis et al., 2021]. There are likely strong links between these methods and what is proposed here.
4 Functional prediction sets
Given a pre-trained model $\hat{f}:X\mapsto F$, which maps from Euclidean space $X\in\mathbb{R}^{m}$ (space of PDE inputs) to a space of functions $F\in\mathcal{F}$ (PDE solutions), discretised on a fine grid $F_{1}=[F_{1}(y_{1}),\dots,F_{1}(y_{l})]\in\mathbb{R}^{l}$, we can compute the error of $\hat{f}$’s error on some unseen data $\{(X_{1},F_{1}),\dots,(X_{n},F_{n})\}$ with $e_{i}=F_{i}-\hat{f}(X_{i})$. We may then reduce the dimension of $e_{i}$ using an SVD: $e=U\Sigma V^{\top}$, where the singular values indicate the variance contribution of each eigenvector to the variance of $e_{i}$. We may truncate the dimensions of the data, capturing some high (often $99\%$) of the overall variance, and projecting the data $e_{i}$ onto these remaining dimensions, a low dimension representation of $\hat{f}$’s error: $U_{i}$. Depending on the size of these dimensions, one of the above fitting methods may be applied to find $\mathcal{Z}^{\alpha}_{p_{Z}}$. We find the ‘rotated hyperrectangle’ method and the Mahalanobis depth scale well to high dimensions, with the convex hull method being tighter for smaller dimensions. $\mathcal{Z}^{\alpha}_{p_{Z}}$ may then be calibrated, giving $\mathbb{P}(U_{n+1}\in\mathcal{Z}^{\alpha})\geq 1-\alpha$. Mapping $\mathcal{Z}^{\alpha}_{p_{Z}}$ through the linear transformation back to $\mathbb{R}^{l}$ is straightforward for the underlying zonotope. The probabilistic component $\alpha$ is also straightforward: each zonotope level-set may be propagated independently $g(\mathcal{Z}^{\alpha}_{p_{Z}})$ for any function $g$, i.e. perform one transformation for each prediction set, with the $\alpha$-guaranteed being retained by the set. Note we may perform this since we compute a multivariate prediction set, this would not be the case had each dimension been fit independently.
These nested sets can be added to the prediction of $\hat{f}$, yielding a functional prediction set
$\mathbb{C}^{\alpha}=\hat{f}(X_{n+1})+U\Sigma\mathcal{Z}^{\alpha}_{p_{Z}}.$
@@ 4.1 Bounding truncation error
To maintain rigorous coverage guarantees, we bound the uncertainty in the truncated dimensions by constructing a hyperrectangle $E$ that encloses the projection of all calibration errors onto the discarded SVD modes. Specifically, we compute the element-wise minimum and maximum of the projected errors in the truncated space, forming a bounding box. Upon taking the Cartesian product $\mathcal{R}^{\alpha}=\mathcal{Z}^{\alpha}_{p_{Z}}\times E$, one obtains a zonotope in high dimensions which contracts in the important directions as $\alpha$ varies, but remains constant in the truncated dimensions. The $\alpha$-guarantee remains unchanged due to this operation because: 1) the number of data points remains unchanged (only their dimension), and 2) the data is totally bounded (and remains so as $\alpha$ changes) in the extra dimensions. Slightly more formally, the indicator function for a Cartesian product is
$\mathbb{I}_{\mathcal{Z}\times E}(x_{1},x_{2})=\mathbb{I}_{\mathcal{Z}}(x_{1})\mathbb{I}_{E}(x_{2}),$
and since $E$ is a bounding box for the calibration data, $\mathbb{I}_{E}$ always returns ‘true’ during calibration, and thus would not change the scoring in equation 4 had it been included. The truncated dimensions may thus be ignored during calibration. Of course, when empirically testing the coverage of a prediction set constructed this way, both the zonotope membership $\mathbb{I}_{\mathcal{Z}}$ and the bounding box membership $\mathbb{I}_{E}$ must be considered.
Indeed, there is a slight performance-loss due to this (our probabilistic bound becomes looser), however this can be expected to be quite minor, as these truncated modes contribute minimally to the overall variance of $F_{i}$. Our final prediction set becomes
$\mathbb{C}^{\alpha}=\hat{f}(X_{n+1})+\mathcal{R}^{\alpha},$
where $\mathcal{R}^{\alpha}=U\Sigma(\mathcal{Z}^{\alpha}_{p_{Z}}\times E)$ has been projected back.
Cartesian product of zonotopes
The Cartesian product between two zonotopes $\mathcal{Z}_{X}\subset\mathbb{R}^{n}$ and $\mathcal{Z}_{Y}\subset\mathbb{R}^{m}$ is a zonotope $\mathcal{Z}_{X\times Y}\subset\mathbb{R}^{n\times m}$, whose centre and generator are a simple concatenation of the centres and generators of $\mathcal{Z}_{X}$ and $\mathcal{Z}_{Y}$: $c_{X\times Y}=[c_{X}\;c_{Y}]$ and $G_{X\times Y}=[G_{X}\;G_{Y}]$.
4.2 PDE surrogate examples
We use LazySets.jl [Forets and Schilling, 2021] for the set construction, and NeuralOperators.jl [Pal, 2023] and some prior models from literature [Gopakumar et al., 2024a] were used for the base Sci-ML models.
Figure 8: Examples of constructed prediction sets of the FNO for Burger’s equation, predicted on test data set. The method predicts nested zonotopes in the high-dimensional space of the surrogate model’s prediction. This plot shows the axis-projection of these zonotopes. The ground truth is shown in yellow.
@@ 4.2 PDE surrogate examples
To demonstrate the methodology, we construct functional prediction sets on a Fourier neural operator (FNO) for the Burger’s equation, trained on data supplementary data provided from Li et al. [2020], which was built by sampling a numerical PDE solver. After taking an FNO model from literature $1025$ points we used to train the SVD, leaving $500$ for calibrating the prediction sets, and $503$ for validation. Additional information about the training setup and the underlying PDEs are provided in appendix D. The predictions from the FNO are discretise onto a grid of 1024 points. A $32$-dimensional rotated hyperrectangle was fitted on the important directions the error’s SVD, with $p_{Z}$ found using the Mahalanobis depth. The entire calibration process took $75$s on a modern laptop, with most of the computation spent on computing the membership 4 of the calibration data. Once calibrated, computing a prediction is timed at $0.02$s. Four realisations of the training dataset are shown in Figure 8, with the ground truth shown in yellow.
@@
We also note that a 1024 dimensional nested zonotopic set is predicted for the output of the FNO, what’s shown in Figure 8 is the axis projection of this set. If we inspect the individual axis, the dependence of the field can also be seen captured.
Table 1: Comparison of empirical coverage and efficiency for various PDEs. Coverage is the empirical frequency with which the true function lies within the predicted set (empirical validation of Eq. 1). Efficiency is a measured as the volume of the prediction set — smaller values indicate tighter predictions for the same coverage level.
$\alpha=0.1$ Modulation Rotated Zonotope
$\alpha=0.2$ Box (ours) (ours)
PDE (model) Coverages
Burgers (FNO) $89.36$ $91.41$ $88.57$
$78.32$ $83.98$ $79$
Burgers (DeepO) $87.79$ $86.8$ $90.22$
$76.46$ $79.15$ $82.24$
Wave (FNO) $86$ $89.2$ $90$
$78$ $78.4$ $81.6$
Navier Stokes $86.83$ $89.17$ $87.33$
(FNO) $75.83$ $78.67$ $79.67$
Efficiency
Burgers (FNO) $3.956$ $2.927^{-1}$ $\mathbf{1.910^{-1}}$
$1.634$ $2.562^{-1}$ $\mathbf{1.867^{-1}}$
Burgers (DeepO) $2.742^{2}$ $\mathbf{3.79}$ $5.787$
$1.049^{2}$ $\mathbf{2.735}$ $5.386$
Wave (FNO) $3.588^{4}$ $5.656^{-3}$ $\mathbf{5.258^{-3}}$
$3.559^{4}$ $\mathbf{4.569^{-3}}$ $4.744^{-3}$
Navier Stokes $9.251^{4}$ $\mathbf{3.423^{1}}$ $3.529^{1}$
(FNO) $7.040^{4}$ $\mathbf{2.441^{1}}$ $3.282^{1}$
$X^{Y}$ denotes $X\times 10^{Y}$
Table 1 gives a more in depth comparison of our method to the supremum-based method (labelled modulation) from Diquigiovanni et al. [2022], for various PDEs and models. We note that our method performs favorably in terms of the tightness of the predicted multivariate set (labelled efficiency). Standard metrics for efficiency for multivariate conformal prediction are to use the set’s volume [Messoudi et al., 2022]. However, since computing the volume of a high dimensional zonotope is challenging, we elect to use the average 2-dimensional projections. Further details of these benchmarks can be found in appendix C and D, including a comparison with lower dimensional data sets from the Mulan and the UCI repository, using more standard MLPs. We note that several other multivariate conformal prediction method are unable to produce results for these models, as they do not scale to the required dimensions. A comparison with these methods is found the appendix in Table 3 for MLPs. We note that for lower dimensional examples, elliptical set conformal prediction performs best, however our method is still competitive.
5 Discussion
@@
The proposed method is similar to the usual score-based conformal prediction, but where we directly solve for a valid prediction set, instead of taking a level-set of a non-conformity scoring function. Indeed, the membership function of $\mathcal{Z}^{\alpha}_{p_{Z}}$ is the non-conformity score in conformal prediction [Gupta et al., 2022]. If one could find an efficient (perhaps analytical) expression for this membership function, in terms of $\alpha$ and the underlying set-representation, then the calibration procedure would be made much more efficient (corresponding to a simple function evaluation), without the need to check the set-membership of $\mathcal{Z}^{\alpha}_{p_{Z}}$ in 4, which is a costly part of the proposition. This would also additionally greatly simplify the implementation.
@@
We therefore share many of the (dis-)advantages of conformal prediction, including finite-sample guarantees and being model agnostic. Although we show example of regression problems, we believe our method (perhaps with a different set-representation) could apply to multivariate classification problems.
Limitations
@@
Since method requires the same assumption as conformal prediction, namely exchangeability, we therefore suffer similar limitations. The guarantee is lost if the dataset is not exchangeable (for example if it changes over time). We also give provide marginal coverage, rather than the stronger conditional coverage, which is more desirable for surrogate modelling. However, methods for improving conditional coverage estimates [Plassier et al., 2025] could be applicable. Other than the usual limitations from conformal prediction, we additionally require an SVD to be trained, which can require a substantial amount of extra training data, in addition to the extra calibration data required for split conformal prediction.
Finally, we mention that our method would greatly benefit for refined methods for fitting zonotopes to data, to improve the tightness of the fitted set.
Acknowledgements.